OpenStudy (anonymous):

PLEASE HELP ME Yves has bought 420 tropical fish for a museum display. He bought 6 times as many parrotfish as angelfish. How many of each type of fish did he buy? 30 angelfish, 180 parrotfish 60 angelfish, 360 parrotfish 120 angelfish, 200 parrotfish 360 angelfish, 60 parrotfish

OpenStudy (anonymous):

30 angle fish and 180 parrot fish, 30X6=180

OpenStudy (anonymous):

sorry @TheOneAndOnlyKacy but it is b

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

what @llhome

OpenStudy (anonymous):

its B right?

OpenStudy (anonymous):

^^^??

OpenStudy (texaschic101):

a + p = 420 p = 6a a + 6a = 420 7a = 420 a = 420/70 a = 60 a + p = 420 60 + p = 420 p = 420 - 60 p = 360 angel fish (a) = 60 parrot fish (p) = 360

OpenStudy (anonymous):

so B @texaschic101

OpenStudy (anonymous):

Yes it is B sorry about that

OpenStudy (anonymous):

i think I, I'm going to go with B

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no Prob